Annaberry22:

Arezoo youth tickets cost five dollars and adult tickets cost nine dollars for group that purchased extra youth tickets and why adult tickets spend $90 on tickets what is the domain of the relationship

mxddi3:

Welcome to QuestionCove (: Are there some answer choices to go with this?

mxddi3:

Also, we are going to label our youth tickets as `x` and our adult tickets as `y`. Then, if the total was $90, this means that 5x+9y=90 (since it's $5 for a youth ticket, $9 for an adult ticket, and the total was $90). Now, domain is is x-values. So we need to solve for this. if we have 5x+9y=90 we need to move the 9y to the other side. we are then left with 5x=90-9y then from here, how do we isolate the "x" variable?